OpenStudy (anonymous):

How do I find the solution to this system? Step by step, please :) {3x + 3y = 10 {-9x + 9y = -30 Please help!

OpenStudy (anonymous):

You can use cramers rule, matrix row operations or substitution. You have a preference?

OpenStudy (anonymous):

Not entirely. Algebra is my weakest subject, you see. So...no.

OpenStudy (anonymous):

Ok, substitution is the easiest conceptually. The idea is you solve for one variable in one equation then plug it into the other. If you divide the top equation by three you get: \[\frac{ 3x }{ 3 }+\frac{ 3y }{ 3 }=\frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

\[x+y=\frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

RIghteo boss, I'm with you so far.

OpenStudy (anonymous):

solve for x: \[x=\frac{ 10 }{ 3 }-y\]

OpenStudy (anonymous):

okee doki

OpenStudy (anonymous):

now plug your value for x into the other equation. We can first divide by a -9, to make the numbers a little smaller.

OpenStudy (anonymous):

Do I have an x factor?

OpenStudy (anonymous):

\[\frac{ -9x }{-9 }+\frac{ 9y }{ -9 }=\frac{ -30 }{ -9 }\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[x-y=\frac{ 30 }{ 9 }\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now remember \[x=\frac{ 10 }{ 3 }-y\] right

OpenStudy (anonymous):

so substitute: \[(\frac{ 10 }{3 }-y)-y=\frac{ 30 }{ 9 }\]

OpenStudy (anonymous):

you with me?

OpenStudy (anonymous):

yeah, I think so

OpenStudy (anonymous):

then combine terms and solve: \[\frac{ 10 }{3}-2y=\frac{ 30 }{ 9 }\]

OpenStudy (anonymous):

Um, but I'm still unsure of what y is. I am really bad at this stuff. (thank you sooo much for your help by the way)

OpenStudy (anonymous):

Also, how do I find Y?

OpenStudy (anonymous):

Are you sure you copied the problem correctly, with the negatives in the right places?

OpenStudy (anonymous):

Yep. Why? Is there something wrong?

OpenStudy (anonymous):

Yes, you get zero like this, let me double check everything.

OpenStudy (anonymous):

Okee doki.

OpenStudy (anonymous):

I double checked, y has to be zero for that system. Which makes x = 3 1/3

OpenStudy (anonymous):

oh, okee doki!

OpenStudy (anonymous):

@JakeV8 check this for me would ya

OpenStudy (anonymous):

Oh, um, hi!

OpenStudy (anonymous):

lots to read, but checking it now. Elimination method might have been easier, but lemme look...

OpenStudy (anonymous):

Please help, I'm a little confused

OpenStudy (anonymous):

Oh, okee doki!

OpenStudy (anonymous):

you get 18y=0?

OpenStudy (anonymous):

me? or Jake?

OpenStudy (anonymous):

Jake, sorry, I just wanted him to double check, its odd you were given a problem with a zero and an irrational answer.

OpenStudy (anonymous):

It's perfectly alright

OpenStudy (anonymous):

Equation 1: 3x + 3y = 10 Equation 2: -9x + 9y = -30 Multiply equation 1 by 3 all the way across... Eq 1: 9x + 9y = 30 Eq 2: -9x + 9y = -30 Then add both equations: 9x - 9x + 9y + 9y = 30 - 30 Simplify: 18y = 0 Solve for y: y = 0 Substitute y = 0 into either equation to solve for x: Eq 1: 3x + 3y = 10 3x + 3(0) = 10 3x = 10 x = 10/3 Final check: put both x = 10/3 and y = 0 into the other equation: Eq 2: -9x + 9y = -30 -9(10/3) + 9(0) = -30 -30 + 0 = -30 <<<--- TRUE, so we got it right :)

OpenStudy (anonymous):

Righteo!

OpenStudy (anonymous):

Sorry, that's a lot all at once.. I can explain if you have questions about anything. Do you understand how I started it off? I wanted to multiply Eq 1 by something so that when I added the equations, one variable would be eliminated.

OpenStudy (anonymous):

Well done Jake. I think the problem was set up to cancel just like you did.

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

yep, me too. Substitution would work fine too.. I just do whatever seems easiest. I find it harder when they tell you to use a certain method that doesn't seem easiest to me.

OpenStudy (anonymous):

@CyberCierra, does this make sense? Algebra doesn't have to be your favorite or best class, but make sure you're learning enough to get through it :)

OpenStudy (anonymous):

I understand it now! Certainly! If you wouldn't mind, could I bother you for a bit more help?

OpenStudy (anonymous):

Can you post it in a new question? @p505 can continue to help to if desired :)

OpenStudy (anonymous):

Righteo boss! Thank you both!